OpenStudy (anonymous):

(Vo - (C1*kR)/2)*kR = - (C1*R^2)/2

OpenStudy (anonymous):

i want to know C1

OpenStudy (lgbasallote):

(Vo -(C1*kr)/2) = -kR(C1*R^2)/2...then... Vo - (C1*kR) = -kR(C1*R^2) -C1*kR = -kR(C1*R^2) - Vo C1 = C1*R^2 + Vo C1 - C1*R^2 = Vo C1(1 - R^2) = Vo C1 = Vo/(1-R^2)

OpenStudy (anonymous):

why -kR its multiplying at the first step?

OpenStudy (anonymous):

failure